Can someone help me with this equation below??
-0.52z = -0.13

Answer by tutorcecilia(2152)

-0.52z = -0.13
-0.52z/-0.52 = -0.13/-0.52
z=0.25
.
check by plugging (z=0.25) back into the original equation and solve
-0.52z = -0.13
-0.52(0.25)z = -0.13
-0.13=-0.13 [checks out]